Problem 3 Review Part A has 7.00x10-22 J more potential energy than a water molecule aligned with the field. The dipole What is the strength of the electric field? Express your answer with the appropriate units. alue N/C Submit K Return to Assignment Provide Feedback
Answer #1

Given,

The change in electric potential energy, U2 - U1 = 7 * 10-22 J

The dipole moment, p = 6.2 * 10-30 Cm

where, U1 is the potential energy perpendicular to E

U2 is the potential energy aligned with the field

We know,

U2 - U1 = p * E

7 * 10-22 =6.2 * 10-30 * E

E = 1.129 * 108 N/C